OpenStudy (anonymous):

Two cites are 64.6 km apart. Vehicle A leaves one city at the same vehicle B leaves the other. find the speed of vehicle A if it travels 6.9 km/h faster than vehicle B if they pass in 15.9 minutes.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

speed = distance / time distance = speed * time Use the second equation to write an equation for each vehicle using the info in the table above. The solve the system of equations. You are looking for x, the speed of vehicle A.

OpenStudy (anonymous):

how do you write the first equation?

OpenStudy (mathstudent55):

distance = speed * time x = (y + 6.9) * (15.9/60)

OpenStudy (anonymous):

now, Do I have to use the result of the first equation to find the second one?

OpenStudy (mathstudent55):

No. Now you use the information about vehicle B to write another equation similar to the one I wrote above. Use again distance = speed * time to write a second equation using the distance, speed and time of vehicle B. Then when you have both equations written, solve the system of equations.

OpenStudy (mathstudent55):

Eq. for vehicle A: x = (y + 6.9) * (15.9/60) Eq. for vehicle B: 64.6 - x = y 8 (15.9/60) Now solve the two equations above simultaneously. You are looking for x.

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

wlcm